A firm has a debt-to-asset ratio of 30% (based on the market value of assets). The firm’s bondholders require a return of 4.50%, and the equity holders require a return of 9%. The firm’s marginal tax-rate is 21%. Estimate the firm’s Weighted-Average-Cost-of-Capital (WACC).
Answer #1

Formula for WACC:

WACC is calculated by multiplying the cost of each capital source (debt and equity) by its relevant weight,and then adding the products together to determine the WACC value.Mathematically

WACC: E/V* Re+D/V*Rd*(1-Tc)

Where ,

E=Market value of the firms equity

D=Market Value of firms debt

V=E+D

Re=Cost of Equity

Rd=Cost of Debt

Tc=Corporate Tax Rate

WACC=9/13.5*0.09+4.5/13.5*4.5(1-0.21)

=0.0599999+1.185=1.245